two bodies begin to fall freely from the same height but the second falls T seconds after the first the time(after which the first body begin to fall) when the distance between the bodies equal to L is
options
A.T/2         B.T/2+L/gT           C.L/gT          D.T+2L/gT

Grade:11

2 Answers

Arun
5 years ago
We have S = 1/2 gT^2 and s = 1/2 g(T - t)^2; find T when S - s = L = 1/2 g (T^2 - (T^2 - 2Tt + t^2)) 

2L/g = 2Tt - t^2 = (2T - t)t 
2L/gt + t = 2T; T = L/gt + t/2 
 
Hence option B is correct
Apurva Sharma
5 years ago
The distance covered by first ball at time t be S1=1/2 gt^2At same time distance covered by second ball be S2= 1/2g (t-T)^2The distance between two balls L= S1 - S2S1-S2 = 1/2g [t^2 -(t-T)^2] =LL=1/2g (t^2 - t^2 + T^2 -2tT)L=1/2gT (T+2t)2L/gT = T+2tThus,answe will be T+2L/gT